OpenStudy (anonymous):

In solving the proportion x-3/4=1/4 , which of the following would be your first step? 4(x - 3) = 1 1(x - 3) = 4(4) 4(x - 3) = 4(1) 1(x - 3) = 4

OpenStudy (mertsj):

Cross multiply. Which choice is the cross products?

OpenStudy (mertsj):

\[\frac{x-3}{4}=\frac{1}{4}\]

OpenStudy (anonymous):

oh thx @Mertsj

OpenStudy (anonymous):

is it the first one or the third one?? @Mertsj

OpenStudy (mertsj):

All depends on what 1 x 4 is